# Multimodal RewardBench 2: Evaluating Omni Reward Models for Interleaved Text and Image

### v1

@@ 1 Introduction
  • We study task-specific metrics (e.g., VQAScore (Lin et al., 2024)) and reward models trained on human preferences (e.g., ImageReward (Xu et al., 2023), UnifiedReward (Wang et al., 2025)), and find that they are no longer reliable on the challenging prompts and frontier models in MMRB2. For instance, VQAScore (with Qwen2.5-VL backbone) and ImageReward achieve 58.3% and 54.0% on text-to-image evaluation, respectively, well below MLLM-as-a-judge baselines such as Qwen3-VL-32B (64.1%) and Gemini 3 Pro (74.4%). While human preference training improves performance over heuristic metrics, these models still fall short of frontier MLLMs.
@@ 2 Related Work
Reward modeling for visual generation. Building on RLHF, reward modeling has been extended beyond text. ImageReward (Xu et al., 2023), HPSv2 and v3 (Wu et al., 2023b; Ma et al., 2025), PickScore (Kirstain et al., 2023) learn human preferences for text-to-image generation, improving correlation with human judgments and guiding diffusion models beyond CLIP-based proxies. For image editing, EditScore (Luo et al., 2025) and EditReward (Wu et al., 2025) adopt similar preference-learning frameworks. Unified approaches aim for cross-task generalization: Wang et al. (2025) train a single multimodal reward across image, video, and understanding tasks. Despite progress, most multimodal RMs remain task-specific and lack a unified, stress-testing evaluation.
Evaluating reward models. Benchmarking reward models has become an active research direction. In the text domain, RewardBench and RewardBench 2 (Lambert et al., 2025; Malik et al., 2025) systematically compare LLM reward functions across diverse axes (e.g., instruction following, reasoning, safety). VL-RewardBench (Li et al., 2025) and Multimodal RewardBench (Yasunaga et al., 2025) assess reward models for multimodal LLM. However, they only focus on image-text-to-text tasks. For image generation, researchers develope automatic evaluation metrics for text-to-image generation. CLIPScore (Hessel et al., 2021) offers a reference-free image–text similarity measure that correlates with human judgments but often misses compositional errors; TIFA (Hu et al., 2023), VQAScore (Lin et al., 2024) address this by probing alignment via VQA, improving robustness on compositional cases. The human annotations collected in these works are often used as reward model evaluations. Most existing reward model evaluations focus either on text or text-to-image generation, offering little insight into interleaved text and image. To bridge this gap, Multimodal RewardBench 2 (MMRB2) provides a unified and challenging framework for assess reward modeling for omni models.
@@ 4.2 Performance of supervised reward models
Judge
Text to
Image
Image
Editing*
Multimodal
Reasoning*
@@
Setup. Unless otherwise noted, we adopt the default metaprompt provided by each official library. For text-to-image, we consider two families of evaluators. The first is CLIP-based (Radford et al., 2021), including CLIPScore (Hessel et al., 2021) and its preference-trained variants ImageReward (Xu et al., 2023), HPSv2 (Wu et al., 2023a), and PickScore (Kirstain et al., 2023). The second family is based on Qwen2.5-VL-7B (Bai et al., 2025). We evaluate VQAScore (Lin et al., 2024), which scores generated images using model logits, as well as the preference-trained reward models HPSv3 (Ma et al., 2025) and UnifiedReward (Wang et al., 2025). We evaluate all of the above models on the MMRB2 text-to-image task. Qwen2.5-VL-7B has also been used as the backbone for reward models on other tasks, including EditReward (Wu et al., 2025) for image editing and UnifiedReward (Wang et al., 2025) for multimodal understanding. Because EditReward is trained only on single-image editing, and UnifiedReward is trained on single-image image-to-text tasks, we evaluate them on the corresponding single-image subsets of MMRB2 to ensure a fair comparison among evaluators. Table 3 summarizes the results.

### v3

@@ 1 Introduction
  • We study task-specific metrics (e.g., VQAScore (Lin et al., 2024)) and reward models trained on human preferences (e.g., ImageReward (Xu et al., 2023), UnifiedReward (Wang et al., 2025b)), and find that they are no longer reliable on the challenging prompts and frontier models in MMRB2. For instance, VQAScore (with Qwen2.5-VL backbone) and ImageReward achieve 58.3% and 54.0% on text-to-image evaluation, respectively, well below MLLM-as-a-judge baselines such as Qwen3-VL-32B (64.1%) and Gemini 3 Pro (74.4%). While human preference training improves performance over heuristic metrics, these models still fall short of frontier MLLMs.
@@ 2 Related Work
Reward modeling for visual generation. Building on RLHF, reward modeling has been extended beyond text. ImageReward (Xu et al., 2023), HPSv2 and v3 (Wu et al., 2023b; Ma et al., 2025), PickScore (Kirstain et al., 2023) learn human preferences for text-to-image generation, improving correlation with human judgments and guiding diffusion models beyond CLIP-based proxies. For image editing, EditScore (Luo et al., 2025) and EditReward (Wu et al., 2025) adopt similar preference-learning frameworks. Unified approaches aim for cross-task generalization: Wang et al. (2025b) train a single multimodal reward across image, video, and understanding tasks. Despite progress, most multimodal RMs remain task-specific and lack a unified, stress-testing evaluation.
Evaluating reward models. Benchmarking reward models has become an active research direction. In the text domain, RewardBench and RewardBench 2 (Lambert et al., 2025; Malik et al., 2025) systematically compare LLM reward functions across diverse axes (e.g., instruction following, reasoning, safety). VL-RewardBench (Li et al., 2025) and Multimodal RewardBench (Yasunaga et al., 2025) assess reward models for multimodal LLM. Llava-Critic series (Xiong et al., 2025; Wang et al., 2025a) focus on developing reward models for these reward benchmarks. EvalPlanner (Saha et al., 2025) and J1 (Whitehouse et al., 2025) further improve reward modeling by incentivizing test-time scaling in LLM-as-a-judge. However, existing benchmarks and judge training efforts still largely focus on image-text-to-text tasks. For image generation, researchers develop automatic evaluation metrics for text-to-image generation. CLIPScore (Hessel et al., 2021) offers a reference-free image–text similarity measure that correlates with human judgments but often misses compositional errors; TIFA (Hu et al., 2023), DSG (Cho et al., 2024), and VQAScore (Lin et al., 2024) address this by probing alignment via VQA, improving robustness on compositional cases. OmniVerifier (Zhang et al., 2025) further investigate on training better visual-outcome verification mdiirhtdkltrnlldrnvhbccugtflukgivodels. The human annotations collected in these works are often used as reward model evaluations. Most existing reward model evaluations focus either on text or text-to-image generation, offering little insight into interleaved text and image. To bridge this gap, Multimodal RewardBench 2 (MMRB2) provides a unified and challenging framework for assess reward modeling for omni models.
@@ 4.2 Performance of supervised reward models
Judge Text to Image Image Editing* Multimodal Reasoning*
@@
Setup. Unless otherwise noted, we adopt the default metaprompt provided by each official library. For text-to-image, we consider two families of evaluators. The first is CLIP-based (Radford et al., 2021), including CLIPScore (Hessel et al., 2021) and its preference-trained variants ImageReward (Xu et al., 2023), HPSv2 (Wu et al., 2023a), and PickScore (Kirstain et al., 2023). The second family is based on Qwen2.5-VL-7B (Bai et al., 2025). We evaluate VQAScore (Lin et al., 2024), which scores generated images using model logits, as well as the preference-trained reward models HPSv3 (Ma et al., 2025) and UnifiedReward (Wang et al., 2025b). We evaluate all of the above models on the MMRB2 text-to-image task. Qwen2.5-VL-7B has also been used as the backbone for reward models on other tasks, including EditReward (Wu et al., 2025) for image editing and UnifiedReward (Wang et al., 2025b) for multimodal understanding. Because EditReward is trained only on single-image editing, and UnifiedReward is trained on single-image image-to-text tasks, we evaluate them on the corresponding single-image subsets of MMRB2 to ensure a fair comparison among evaluators. Table 3 summarizes the results.